The following financial information is available for Marigold Corporation.

(in millions)

2022

2021

Average common stockholders’ equity

$3,000 $3,200

Dividends declared for common stockholders

370 714

Dividends declared for preferred stockholders

40 40

Net income

625 680


Calculate the payout ratio and return on common stockholders’ equity for 2022 and 2021. (Round answers to 1 decimal place, e.g. 12.5%)

2022

2021

Payout ratio

enter payout ratio in percentages % enter payout ratio in percentages %

Return on common stockholders’ equity

enter return on common stockholders’ equity ratio in percentages % enter return on common stockholders’ equity ratio in percentages %
Answer #1

Payout ratio = dividends declared to common stockholders/Net income.

Return on common stockholders equity = (Net income - Preferred dividends)/Average common stockholders equity


20222021
Payout ratio

= 370/625

= 59.2%

= 714/680

= 105%

Return on common stockholders equity

= (625-40)/3000

= 19.5%

= (680-40)/3200

= 20%


answered by: helium